OpenStudy (anonymous):

Tim has some money to buy deepavali greetings cards. He has $6 left after he has bought 5 cards. However, he will need $3 more if he wants to buy 6 cards. How much money does he have?

OpenStudy (nubeer):

she is left with 6$ and need 3$ more to buy a card .. it makes 9$ per card.. she has bought 5 cards. so price of 5 cards = 9*5 = 45$ plus she is left with 6$ so total money 45 + 6 = 51$

OpenStudy (shubhamsrg):

let money he had be x let cost of 1 card = y now, x-5y = 6 and x+3 = 6y solve both..

OpenStudy (shubhamsrg):

she* :P